5 years ago, Barton Industries issued 25-year noncallable, semiannual bonds with a $1,000 face value and a 9% coupon, semiannual payment ($45 payment every 6 months). The bonds currently sell for $896.87. If the firm's marginal tax rate is 25%, what is the firm's after-tax cost of debt? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Solution: After tax cost of debt = Before tax cost of debt x (100 - Tax Rate) Or After tax cost of debt = Bond Yield to matur
Answer #2

SOLUTION :


Using online calculator for YTM ,  ref : https://dqydj.com/bond-yield-to-maturity-calculator/ :


Key -in the following :


Current market price ($) = 896.87


Bond Face Value ($) = 1000


Years to maturity (25 - 5) = 20 


Annual coupon rate (%) = 9.0


Frequency of coupon = 2 (being semi-annual) 



And see the result : YTM = Pre tax cost of debt = 10.22 %  


So, 


After tax cost of debt = 10.22( 1 - 0.25) = 7.665 = 7.66 % (ANSWER)

answered by: Tulsiram Garg